4. Assume that the equilibrium in the loanable funds market is at interest rate of 1.25% and quantity of funds at $20 billion. Suppose the current government deficit is zero so government is not borrowing any money.

a) Suppose now government increases spending by $2 billion and finances it entirely by borrowing. This deficit increases equilibrium interest rate to 2% and equilibrium quantity of funds to $21.5. Show the changes on the graph.

b) What happens to private investment (I) and total savings as a result of increase in the government spending?

c) Assuming total crowding out, that is, no change in aggregate demand, by how much does consumption (C) change?
Answer #1

(a)

Increase in government spending will increase government borrowing, increasing the demand for loanable funds. The loanable funds demand curve will shift to right, increasing interest rate and increasing quantity of loanable funds.

In following graph, D0 and S0 are initial demand and supply curves of loanable funds, intersecting at point A with initial interest rate r0 (= 1.25%) and quantity of loanable funds Q0 (= $20 billion). As demand rises, D0 shifts right to D1 (where D1 - D0 = $2 billion), intersecting S0 at higher interest rate r1 (= 2%) and higher quantity of loanable funds Q1 (= $21.5 billion).

80 YDO DI 81 8

(b)

Increase in investment and saving = Q1 - Q0 = $21.5 billion - $20 billion = $1.5 billion

(c)

Y = C + I

\DeltaY = \DeltaC + \DeltaI

0 = \DeltaC + $1.5 billion

\DeltaC = - $1.5 billion (decrease by $1.5 billion)